# Engagement Maximization

### v5

@@ 3.3 The agent-optimal benchmark and belief polarization
$V^{B}(q)>\hat{u}(q)$implies $\bar{J}(q)>0$. Moreover, $\forall\pi$ that implements $V^{B}(\bar{q}_{0})$, $\mathbb{E}^{\pi}[V^{B}(q)-\hat{\mu}(q)]=0$.
@@
This lemma highlights two points. First, if at some $q$ the agent would choose to acquire information in the agent-optimal benchmark, the principal can profit by providing the agent with information (although that information will not in general be the information the agent would have chosen to acquire). Second, at any belief the agent might hold when stopping in the agent-optimal benchmark, the agent must have no benefit from continuing to acquire information.
The static rational inattention problem that characterizes stopping beliefs in the agent-optimal benchmark is essentially identical to the one that characterizes optimal policy in our principal-agent framework, except that the UPS information cost is $\frac{\bar{\kappa}}{\chi}H(q)$ in the benchmark problem and $\frac{\bar{\kappa}}{\chi}H(q)-\lambda G(q)$ in the principal agent case. The similarity in structure between these two problems will allow us to highlight the implications of engagement maximization.
Let $Q^{i}(\bar{q}_{0})\subseteq\mathcal{P}(X)$ be the union of the support of all optimal stopping beliefs in the benchmark ($i=a$) and principal-agent ($i=p$) models. Let $\mathrm{Conv}\ Q^{i}(\bar{q}_{0})$ denote the convex hull of $Q^{i}(\bar{q}_{0})$. The following proposition demonstrates that the beliefs the agent will hold when stopping after using the principal’s platform are more extreme than the beliefs the agent would choose to acquire in the benchmark model. This result follows from the observation that in both models, stopping beliefs are characterized by the solution to a static rational inattention problem, with a lower information cost in the principal-agent case than in the benchmark case. There are two possible exceptions to this result. One is the case in which the principal cannot induce any engagement ($\bar{J}(q_{0})=0$ in Proposition 1), in which case there is a principal-optimal policy in which no information is acquired, $\bar{q}_{0}\in Q^{p}(\bar{q}_{0})$. The second is the case in which the agent’s beliefs lie on the boundary of the simplex, in which case there may not be “room” for beliefs to become more extreme.
@@
Assume that $\bar{J}(q_{0})>0$ and that $Q^{a}(\bar{q}_{0})$ is a subset of the relative interior of $\mathcal{P}(X)$. Then for all $q\in Q^{p}(\bar{q}_{0}),q\not\in\ \mathrm{Conv}\ Q^{a}(\bar{q}_{0})$.
@@
Leading Case: $H$=$G$.
When the principal benefits from all forms of engagement by the agent (our interpretation of the $H=G$ case), it is never worthwhile for the principal to provide the agent with decision-irrelevant information. Instead, the principal induces the agent to acquire more decision-relevant information than the agent would choose to acquire on her own.
To illustrate this, we suppose that the state space has a product structure, $X=X_{1}\times X_{2}$, and that the agent’s decision problem depends only on $x_{1}\in X_{1}$. We further suppose that $x_{1}$ and $x_{2}$ are independent under the prior, so that an agent who receives signals that depend only on $x_{1}$ will not update about $x_{2}$.
The following result shows that this is what happens in the principal-agent problem– the principal will send the agent information about $x_{1}$ only. This occurs even though the principal would benefit from the agent learning about $x_{2}$ as well. Intuitively, signals about $x_{1}$ and $x_{2}$ that benefit the principal equally take an equal amount of time for the agent to process (because $H=G$), but only the former has a utility benefit for the agent. A principal who was sending signals about $x_{2}$ could switch to sending signals about $x_{1}$, relax the agent’s participation constraint, and then profit by adjusting the signals so that they induced even more extreme beliefs about $x_{1}$ for the agent.
@@
Let $x=(x_{1},x_{2})$, where $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ and $x_{1},x_{2}$ are independently distributed according to $\bar{q}_{0}$. Suppose (i) $\hat{u}(x)$ depends on $x_{1}$ only and (ii) $G=H$, then (R) is solved by $\pi$ that reveals $x_{1}$ only, i.e. for all $q\in\text{Supp}(\pi)$ and all functions $f:X_{2}\rightarrow\mathcal{R}$, $\mathbb{E}^{q}[f(x_{2})]=\mathbb{E}^{\bar{q}_{0}}[f(x_{2})]$.
@@
Leading Case: Decision-Irrelevant $G$
Let us now consider a different leading case, in which the information that benefits the principal is irrelevant to the agent’s decision. Maintaining the assumption of a product structure for the state space and an independent prior, we now suppose that $G$ depends only on $x_{2}$ (recall that $\hat{u}$ depends only on $x_{1}$. To derive a sharp result, we will assume that the $H$ function is the negative of Shannon’s entropy. [^15] We conjecture that the property required for the result is the $R$-mononticity property discussed in Hébert and La’O (2023).
@@
Let $x=(x_{1},x_{2})$, where $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ and $x_{1},x_{2}$ are independently distributed according to $\bar{q}_{0}$. Suppose (i) $\hat{u}(x)$ depends on $x_{1}$ only, (ii) $G$ depends on $x_{2}$ only and (iii) $H$ is the negative of Shannon’s entropy, then (R) is solved by $\pi=\pi^{*}_{1}\otimes\pi_{2}$, where $\pi^{*}_{1}\in\mathcal{P}(\mathcal{P}(X_{1}))$ is optimal for the agent and $\pi_{2}\in\mathcal{P}(\mathcal{P}(X_{2}))$.
@@
When the principal cares only about providing decision-irrelevant information and the rate of information acquisition is bounded by Shannon’s entropy, the principal will provide a mix of decision-relevant and decision-irrelevant information. The decision-relevant information will be exactly the information that the agent would have chosen for herself, but the agent will be unable to avoid learning the decision-irrelevant information that is “mixed in” with the decision-relevant information.
3.4 Implications of Engagement Maximization
In this subsection, we consider several implications of our results thus far.
Implication 1: Agent Welfare Minimization.
$\mathbb{E}[\hat{u}(q_{\tau^{*}})-\bar{\kappa}\tau^{*}|\mathcal{F}_{0}^{*}]=\hat{u}(\bar{q}_{0}).$
Strikingly, the agent is no better off receiving information from an engagement-maximizing principal than if she could not receive any information at all. The principal extracts the full surplus generated by the ability to produce information and apply it in the agent’s decision problem, despite the agent’s ability to choose when to stop and act. The principal, in maximizing the engagement of the agent, minimizes the agent’s welfare subject to a participation constraint.
A different way of viewing this result is as showing that paying the principal for the information using attention as opposed to a monetary payment is inefficient. If the principal could commit to a signal structure in exchange for a payment from the agent (the transferable utility case), he would of course provide the optimal signal structure from the agent’s perspective, as this would maximize revenue. Our model assumes that the principal is instead paid via attention, which distorts the principal’s incentives for information provision.
In practice, there are likely good reasons why platforms such as Facebook and Twitter choose to make their services free to users. Providers that charge users in exchange for information (the New York Times, for example) generally have far fewer users than freely accessible, advertising-supported platforms. Our baseline model does not speak to this trade-off, but instead offers an explanation for why the information available on free-to-use platforms differs in nature from the information available from paid information providers.
Implication 2: Extreme Beliefs.
This result is described in Proposition 3 and is illustrated in Figure 2 below. Extreme beliefs are a natural consequence of engagement maximization. By forcing the agent to ultimately acquire more information than she would choose for herself (the extreme beliefs), the principal can simultaneously delay the agent’s stopping decision while providing information the agent is willing to attend to. Interpreted in the context of an internet platform and a user, our extreme beliefs result implies that the platform should provide in-depth content on a narrower set of topics than would be preferred by the user. We can interpret this loosely as encouraging the user to go down “rabbit holes” that lead to extreme beliefs. In the teacher-student context, extreme beliefs have a more benign interpretation: the teacher induces test-motivated students to learn more than they would choose to learn on their own.
@@ 3.5 Illustrative Examples
We illustrate our results thus far in a simple example. An agent faces a choice between two actions, $A=\{l,r\}$. The payoffs from the actions are uncertain and depend on the state of the world $x\in X=\{L,R\}$. The agent assigns equal prior probability to both states ($\bar{q}_{0}$ is uniform). The agent gets utility one when the chosen action matches the state ($u_{l,L}=u_{r,R}=1$) and utility negative one otherwise ($u_{l,R}=u_{r,L}=-1$). The agent is impatient and pays a constant cost of delay of two utils per unit of time ($\bar{\kappa}=2$). We assume that the information processing constraint $H$ is the negative of Shannon’s entropy, and will begin by assuming that the engagement measurement function $G$ is also proportional to Shannon’s entropy.
To build intuition, we will restrict attention to the following type of dilution signals (which is without loss, as this class includes the optimal signal structure). Each signal has two possible realizations $\{s_{l},s_{r}\}$, which arrives at Poisson rates that depend on the state (symmetrically):
L R
$s_{l}$ $\alpha\cdot p$ $\alpha\cdot(1-p)$
$s_{r}$ $\alpha\cdot(1-p)$ $\alpha\cdot p$
The parameters $\alpha\in\mathbb{R}_{+}$ and $p\in[\frac{1}{2},1]$ represents the arrival frequency and precision of the signal, respectively. In the context of an internet platform sequentially presenting news articles, tweets, or other kinds of content to a user, we will think the platform as presenting content to the user at a constant rate. The parameter $\alpha$ represents the likelihood that the content presented over an instant time is relevant to the decision facing the user. The parameter $p$ represents the depth or informativeness of the content conditional on its relevance.
The information processing constraint for this class of signals requires that $\alpha$ and $p$ must satisfy
$\alpha\cdot\left(H^{S}\left(\frac{1}{2}\right)-H^{S}(p)\right)\leq 1,$
where $H^{S}(\cdot)$ is Shannon’s entropy. Note that $H^{S}\left(\frac{1}{2}\right)-H^{S}(p)$ is an increasing and convex function of $p$. Intuitively, both more frequent and more precise signals require processing “more information.” However, there is a trade-off between the frequency and the precision of signals.
In the context of an internet platform sequentially presenting content to a user, we interpret this constraint as a breadth-vs.-depth trade-off. A news feed can initially present articles on a broad range of potentially relevant topics, increasing the likelihood that one is relevant to the user. Alternatively, a news feed can concentrate its initial coverage on a narrower range of topics, increasing its informativeness (conditional on relevance) but risking the possibility that none of the initial articles are relevant for the user. The former corresponds to a high $\alpha$, low $p$ strategy, the latter to a low $\alpha$, high $p$ strategy. The platform can choose between these strategies, but cannot present in-depth information about a large range of topics in a short period of time and expect the user to the process all of the information.
Consider the signal with $p=\frac{e}{1+e}$ and $\alpha=\frac{1}{H^{S}(\frac{1}{2})-H^{S}(p)}.$ This signal is feasible and the capacity constraint is exactly binding. Suppose the signal arrives: it is straightforward to show that the agent should stop and follow the recommendation of the signal— choosing $l$ ($r$) upon the arrival of $s_{l}$ ($s_{r}$). The agent’s expected utility at time zero under this strategy is
$V^{B}=\underbrace{\frac{e-1}{1+e}}_{\text{Utility from Actions}}-\underbrace{2\alpha^{-1}}_{\text{Expected Cost of Delay}}\approx 0.24.$
We now explain why the agent is willing to wait for the arrival of the signal. If the agent has not observed either signal by time $t$, her beliefs will remain equal to her prior (and hence assign equal probability to each state). As a result, if the agent were to stop and act before observing either of the Poisson signals, her expected payoff would be zero. Because the problem is stationary, $V^{B}$ is also the agent’s expected utility at any time $t>0$, conditional on not having received either Poisson signal yet; it follows that the agent will strictly prefer waiting for the arrival of one of these signals to stopping. In fact, the results of Hébert and Woodford (2023) show that this signal structure and stopping strategy is optimal for the agent, both within the class of signals considered in this example and within the larger class of signals satisfying the mutual information bound. This strategy also yields a positive payoff for the principal. Because we have assumed that the measure of engagement is the negative of Shannon’s entropy, the expected engagement in this case is $\alpha^{-1}$ (which is also the expected decision time).
However, it is possible for the principal to do better. If the prinicpal exactly satisfies the agent’s information processing constraint (which will be optimal), within this class of Poisson signals, maximizing engagement requires increasing the expected decision time. This is only possible if the principal offers signals that are less frequent but more informative. Such signals are less desirable for the agent than the agent-optimal signals (by definition); but as established above, the agent strictly benefits from receiving the agent-optimal signals, and therefore might be willing receive somewhat less desirable signals as well.
Specifically, consider the signal with $\tilde{p}=\frac{4e}{1+4e}$ and $\tilde{\alpha}=\frac{1}{H^{S}(\frac{1}{2})-H^{S}(\tilde{p})}$. Evidently, the signal is more precise conditional on arrival ($\tilde{p}>p$) but arrives less frequently ($\tilde{\alpha}<\alpha$). The signal exactly satisfies the information processing constraint. If the agent still stops and follows the recommendation of the signal upon signal arrival, her expected utility will be
$\tilde{V}^{B}=\underbrace{\frac{4e-1}{1+4e}}_{\text{Utility from Actions}}-\underbrace{2\tilde{\alpha}^{-1}}_{\text{Expected Cost of Delay}}\approx 0.02>0.$
From the agent’s perspective, waiting for this signal is only slightly better than stopping– but it is strictly better, and by the same stationarity arguments given above, the agent will be willing to wait.
Using these less frequent, more informative signals allows the principal to increase expected engagement ($\tilde{\alpha}^{-1}>\alpha^{-1}$) at the agent’s expense. It is straightforward in this example to see that using an even less frequent, even more informative signal would allow the principal to drive the agent to her indifference point while further increasing expected engagement. Note that utility is not transferable in this example (or in our general model); thus, the necessity of welfare minimization under the principal’s optimal strategy is not immediate from the setup of the model, and in fact will not hold when the principal is restricted to a subset of strategies (see Section 5).
This example illustrates the forces behind our main result. By assumption, the principal benefits from sending the agent signals that the agent will attend to, as it is assumed that this is what generates revenue for the principal. The agent perceives delay as costly, and will only continue to attend to the principal’s signals if she values the information being provided. The agent will in general be willing to receive “too much” information, relative to what she would choose for herself, provided that this still leaves her with some surplus. The principal can take advantage of this willingness by sending signals that always result in the agent either learning too much (generating what we call “extreme beliefs”) or leaving beliefs unchanged. These signals generate larger expected stopping times, and thus more revenue for the principal. In the context of an internet platform providing content sequentially to a user, it is optimal for the platform to provide a stream of mostly-irrelevant pieces of content, a few of which will be highly informative, as opposed to providing a stream of brief and somewhat-relevant content.
@@
We next illustrate Theorem 1 in the context of this example. We will now vary the agent’s prior belief $\bar{q}$, and focus our analysis on the agent’s posteriors when stopping ($q_{\tau}$) as opposed to the compound Poisson signals that generate those posteriors. Figure 1 illustrates the optimal $q_{\tau}$ for every prior belief: When $\bar{q}_{0}\leq\underline{q}$ or $\bar{q}_{0}\geq\bar{q}$, the agent stops immediately no matter what information she will receive. When $\bar{q}_{0}\in(\underline{q},\bar{q})$, the dashed lines illustrate the support of the agent-optimal policy. The lines are flat, meaning that the agent’s stopping beliefs do not change with the prior belief, consistent with the prediction of the standard RI models. [^16] This property is known as the “locally invariant posteriors” (Caplin et al. (2022)). The solid curves illustrate the support of the unique optimal $\pi^{*}$. They are closer to the boundaries zero and one, indicating that the posterior beliefs become more polarized under the principal’s optimal policy.
Figure 1: $\text{Supp}(q_{\tau})$ as a correspondence of $q_{0}$
@@
Figure 2 illustrates the engagement level for every prior belief. The dashed curves represent the agent-optimal policy and the solid curves represent the principal-optimal policy. The principal-optimal policy induces higher engagement level than the agent-optimal policy. However, the engagement level converges to zero when the prior belief goes to the boundary of the continuation region $(\underline{q},\bar{q})$.
Figure 2: Engagement as a function of $q_{0}$
@@
We next adapt example 1 to illustrate our two leading cases, in the context of engaging test-motivated students. Suppose the true state of the world $T$ consists of both a test-relevant and test-irrelevant dimension, $T=T_{1}\times T_{2}=\{L,R\}\times\{0,1\}$. The student and teacher know that the student will be asked a single question, $Q=\{Q_{0}\}$, whose responses are $A=\{l,r\}$, with $Q_{0}(L0\}=Q_{0}(L1)=l$ and $Q_{0}(R0)=Q_{0}(R1)=r$. That is, the $\{L,R\}$ component of the true state is relevant, and the $\{0,1\}$ component is irrelevant. The student (agent) again faces a binary choice $A=\{l,r\}.$ The student’s utility is one if she answers the question correctly and negative one otherwise, and therefore is identical (ignoring the test-irrelevant dimension) to that in example 1. The prior belief is uniform. We continue to assume the agent’s information processing capacity is defined by the negative of Shannon’s entropy, and continue to use the parameters $\bar{\kappa}=2,\chi=1$.
@@
We first study the leading case of $H=G$, and illustrate Proposition 4 with a specific example. By Proposition 1, the teacher’s optimal strategy maximizes
$\mathbb{E}^{\pi}[\hat{u}(q)-(2-\lambda)(H(q)-H(\bar{q}_{0}))]$
for some $\lambda<2.$ The student-optimal benchmark maximizes
$\mathbb{E}^{\pi}[\hat{u}(q)-2(H(q)-H(\bar{q}_{0}))].$
Both optimization problem can be written as:
$\sup_{P(a|t_{1},t_{2})}\sum_{a,t_{1},t_{2}}\frac{1}{4}u(a,t_{1})P(a|t_{1},t_{2})-C\cdot(\sum_{a}\sum_{t_{1},t_{2}}\frac{1}{4}P(a|t_{1},t_{2})\log(P(a|t_{1},t_{2})$
$-\sum_{a}(\sum_{t_{1},t_{2}}\frac{1}{4}P(a|t_{1},t_{2}))\log(\sum_{t_{1},t_{2}}\frac{1}{4}P(a|t_{1},t_{2}))),$
for some positive constant $C$. Note that replacing $P(a|t_{1},t_{2}=0)$ and $P(a|t_{1},t_{2}=1)$ with $\frac{P(a|t_{1},0)+P(a|t_{1},1)}{2}$ (denoted by $P(a|t_{1})$) does not change the positive term while strictly reduces the negative term if $P(a|t_{1},0)$ and $P(a|t_{2},1)$ are not identical. Evidently, w.l.o.g., the optimization problem reduces to
$\sup_{P(a|t_{1})}\sum_{a,t_{1}}\frac{1}{2}u(a,t_{1})P(a|t_{1})-C\cdot(\sum_{a,t_{1}}\frac{1}{2}P(a|t_{1})\log(P(a|t_{1}))$
$-\sum_{a,t_{1}}\frac{1}{2}P(a|t_{1})\log(\sum_{t_{1}}\frac{1}{2}P(a|t_{1}))),$
which is equivalent to a static rational inattention problem. [^17] This equivalence is special to mutual information and is the implication of a more general “compression invariance” property introduced by Caplin et al. (2022); Bloedel and Zhong (2020). As shown in Proposition 4, when there is a test-irrelevant state that is costly to learn about, no information about the state will be acquired either in the student-optimal benchmark or in the principal-agent problem, even though the student learning about state $t_{2}$ enters the teacher’s payoff function.
@@
Now, we turn to the second leading case, in which the teacher only cares about how much the student knows about the state $t_{2}\in T_{2}$ (the test-irrelevant dimension). We assume that the teacher’s payoff is $2\mathbb{E^{\pi}}[|q_{t_{2}}-0.5|]$ if the measure of stopping belief is $\pi$, where $q_{t_{2}}$ denotes the probability of $t_{2}=1$ under the student’s stopping belief. [^18] This objective does not satisfy our differentiability assumptions for $G$ at all points, but the issues arising from non-differentiability will not be relevant for this example. Note that this payoff function is equivalent to the instrumental value of information in a hypothetical binary decision problem where the utility of matching (mismatching) the state $y$ is 1 (-$1$). Proposition 1 can be extended to show that the teacher’s optimal policy maximizes the following auxiliary problem (for some $\lambda>0$):
$\sup_{\pi}\mathbb{E^{\pi}}[2|q_{t_{2}}-0.5|+\lambda\hat{u}(q)-2\lambda(H(q)-H(\bar{q}_{0}))],$
which is equivalent to a hypothetical static RI problem where the teacher has four possible actions, whose utilities are
@@
Per Matêjka et al. (2015), the solution is logit, given by

### v6

@@ 3.2 The agent-optimal benchmark and extreme beliefs
If the agent would choose to acquire information in the agent-optimal benchmark, the principal will induce engagement: for all $q\in\mathcal{P}(X)$, if $V^{B}(q)>\widehat{u}(q)$ then $\bar{J}(q)>0$.
Moreover, if $\pi_{A}$ is a solution to (3.2), then starting (3.2) from any belief in the support of $\pi_{A}$, it is weakly optimal for the agent to stop, $\mathbb{E}^{\pi_{A}}[V^{B}(q)-\hat{u}(q)]=0$.
@@
This lemma highlights two points. First, if the agent would choose to acquire information in the agent-optimal benchmark, the principal can profit by providing the agent with information (although that information will not in general be the information the agent would have chosen to acquire). Second, at any belief the agent might hold when stopping in the agent-optimal benchmark, the agent must have no benefit from continuing to acquire information.
The static rational inattention problem that characterizes stopping beliefs in the agent-optimal benchmark is essentially identical to the one that characterizes optimal policy in our principal-agent framework, except that the UPS information cost is $\frac{\kappa}{\chi}H(q)$ in the benchmark problem and $\frac{\kappa}{\chi}H(q)-\lambda G(q)$ in the principal agent case. The similarity in structure between these two problems will allow us to highlight the implications of engagement maximization.
Let $Q^{i}(\bar{q}_{0})\subseteq\mathcal{P}(X)$ be the union of the support of all optimal stopping beliefs in the benchmark ($i=a$) and principal-agent ($i=p$) models. Let $\mathrm{Conv}\ Q^{i}(\bar{q}_{0})$ denote the convex hull of $Q^{i}(\bar{q}_{0})$. The following proposition demonstrates that the beliefs the agent will hold after engaging with the principal are more extreme than the beliefs the agent would choose to acquire in the benchmark model. This result follows from the observation that in both models, stopping beliefs are characterized by the solution to a static rational inattention problem, with a lower information cost in the principal-agent case than in the benchmark case. There are three possible exceptions to this result. One is the case in which the principal can provide no strict incentive via information ($V^{B}(\bar{q}_{0})=0$), in which case $Q^{p}(\bar{q}_{0})\subset Q^{a}(\bar{q}_{0})$. The second is the case in which the agent-optimal beliefs lie on the boundary of the simplex, in which case there may not be “room” for beliefs to become more extreme. The third is the possibility that the marginal benefit to the principal of additional information acquisition jumps downward exactly at the agent-optimal posterior beliefs, which can be ruled out by assuming $G$ is differentiable.
@@
Assume that $G$ is continuously differentiable, $V^{B}(\bar{q}_{0})>\widehat{u}(\bar{q}_{0})$, and that $Q^{a}(\bar{q}_{0})$ is a subset of the relative interior of $\mathcal{P}(X)$. Then for all $q\in Q^{p}(\bar{q}_{0}),q\not\in\ \mathrm{Conv}\ Q^{a}(\bar{q}_{0})$.
@@
Extreme beliefs are a natural consequence of engagement maximization. By forcing the agent to ultimately acquire more information than she would choose for herself (the extreme beliefs), the principal can simultaneously delay the agent’s stopping decision while providing information the agent is willing to attend to. Interpreted in the context of an internet platform and a user, our extreme beliefs result implies that the platform should provide in-depth content on a narrower set of topics than would be preferred by the user. We can interpret this loosely as encouraging the user to go down “rabbit holes” that lead to extreme beliefs. In the teacher-student context, extreme beliefs have a more benign interpretation: the teacher induces test-motivated students to learn more than they would choose to learn on their own.
3.2.1 Leading cases
˜3 establishes that engagement maximization necessarily leads to more extreme stopping beliefs relative to the agent optimal benchmark. In what follows, we provide a more detailed analysis of the direction towards which engagement maximization biases the beliefs in two salient settings. To help interpretation, we assume that the state space has a product structure, $X=X_{1}\times X_{2}$, and that the agent’s decision problem $u$ depends only on $x_{1}\in X_{1}$. We further suppose that $x_{1}$ and $x_{2}$ are independent under the prior $\bar{q}_{0}$, so that an agent who receives signals that depend only on $x_{1}$ will not update about $x_{2}$. Thirdly, we assume throughout this subsection that $H$ is the negative of Shannon’s entropy $H^{S}$.
All Engagement is Profitable.
When the principal benefits from all forms of engagement by the agent (our interpretation of the $G=H$ case), it is never worthwhile for the principal to provide the agent with decision-irrelevant information. Instead, the principal induces the agent to acquire more decision-relevant information than the agent would choose to acquire on her own.
The following result shows that this is what happens in the principal-agent problem– the principal will send the agent information about $x_{1}$ only. This occurs even though the principal would benefit from the agent learning about $x_{2}$ as well. Intuitively, signals about $x_{1}$ and $x_{2}$ that benefit the principal equally take an equal amount of time for the agent to process (because $G=H$), but only the former has a utility benefit for the agent. A principal who was sending signals about $x_{2}$ could switch to sending signals about $x_{1}$, relax the agent’s participation constraint, [^17] This argument depends on the assumption that sending a signal conditioned only on $x_{1}$ is least costly way to communicate information about $x_{1}$; that $H$ is proportional to Shannon’s entropy ensures this. and then profit by adjusting the signals so that they induced even more extreme beliefs about $x_{1}$ for the agent.
@@ 3.2.1 Leading cases
Suppose $G=H$. Then, (R) is solved by $\pi$ that reveals $x_{1}$ only, i.e. for all $q\in\text{Supp}(\pi)$ the marginals of $q$ and $\bar{q}_{0}$ on $X_{2}$ are identical.
@@
Decision-Irrelevant Engagement is Profitable.
Let us now consider a different leading case, in which the information that benefits the principal is irrelevant to the agent’s decision. Maintaining the assumption of a product structure for the state space, an independent prior, and that $H$ is proportional to Shannon’s entropy, we now suppose that $G$ depends only on $x_{2}$ (recall that $\widehat{u}$ depends only on $x_{1}$).
@@
Suppose $G$ depends on $x_{2}$ only. Then, (R) is solved by $\pi=\pi^{*}_{1}\otimes\pi_{2}$, where $\pi^{*}_{1}\in\mathcal{P}(\mathcal{P}(X_{1}))$ is optimal for the agent and $\pi_{2}\in\mathcal{P}(\mathcal{P}(X_{2}))$.
@@
When the principal cares only about providing decision-irrelevant information, the principal will provide a mix of decision-relevant and decision-irrelevant information. The decision-relevant information will be exactly the information that the agent would have chosen for herself, but the agent will be unable to avoid learning the decision-irrelevant information that is “mixed in” with the decision-relevant information. [^18] We conjecture that the property required for the result is the $R$-mononticity property discussed in hebert2023information.
@@ 3.3 Illustrative Examples
We illustrate our results thus far in a simple example. An agent faces a choice between two actions, $A=\{l,r\}$. The payoffs from the actions are uncertain and depend on the state of the world $x\in X=\{L,R\}$. The agent assigns equal prior probability to both states ($\bar{q}_{0}$ is uniform). The agent gets utility one when the chosen action matches the state ($u_{l,L}=u_{r,R}=1$) and utility negative one otherwise ($u_{l,R}=u_{r,L}=-1$). The agent is impatient and pays a constant cost of delay of two utils per unit of time ($\kappa=2$). We assume that the information processing constraint $H$ is the negative of Shannon’s entropy $H^{S}$, and that the engagement measurement function $G$ is also the negative of Shannon’s entropy. We assume $\chi=1$.
By Theorem 1, the principal’s optimal strategy maximizes
$\mathbb{E}^{\pi}[\widehat{u}(q)-(2-\lambda)(H^{S}(0.5)-H^{S}(q))]$
for some $\lambda\in(0,2)$. Meanwhile, the agent’s optimal strategy maximizes
$\mathbb{E}^{\pi}[\widehat{u}(q)-2(H^{S}(0.5)-H^{S}(q))].$
Both problems are static rational inattention problems. An application of matejka2015rational implies that the agent’s optimal stopping belief has support $\{\frac{e}{e+1},\frac{1}{e+1}\}$ and the principal’s optimal stopping belief has support $\Big\{\frac{e^{2/(2-\lambda)}}{e^{2/(2-\lambda)+1}},\frac{1}{e^{2/(2-\lambda)}+1}\Big\}$. Both probability measures of stopping beliefs can be implemented by dilutions (compound Poisson processes). Since $\lambda\in(0,1)$, it is easy to see that the principal induces more extreme posterior beliefs and a longer wait compared to the agent-optimal benchmark.
@@
Next, we vary the agent’s prior belief $\bar{q}$, and focus our analysis on the agent’s posteriors. Figure 1 illustrates the optimal $q_{\tau}$ for every prior belief: When $\bar{q}_{0}\leq\underline{q}$ or $\bar{q}_{0}\geq\bar{q}$, the agent stops immediately no matter what information she will receive. When $\bar{q}_{0}\in(\underline{q},\bar{q})$, the dashed lines illustrate the support of the agent-optimal policy. The lines are flat, meaning that the agent’s stopping beliefs do not change with the prior belief, consistent with the prediction of the standard RI models. [^19] This property is known as the “locally invariant posteriors” (caplin2017shannon). The solid curves illustrate the support of the unique optimal $\pi^{*}$. They are closer to the boundaries zero and one, indicating that the posterior beliefs become more extreme under the principal’s optimal policy.
[…]
Figure 1: $\text{Supp}(q_{\tau})$ as a correspondence of $\bar{q}_{0}$
@@
Figure 2 illustrates the engagement level $\mathbb{E}^{P}[G(q_{\tau})-G(\bar{q}_{0})]$ for every prior belief. The dashed curves represent the agent-optimal policy and the solid curves represent the principal-optimal policy. The principal-optimal policy induces higher engagement level than the agent-optimal policy. However, the engagement level converges to zero when the prior belief goes to the boundary of the continuation region $(\underline{q},\bar{q})$.
[…]
Figure 2: Engagement as a function of $\bar{q}_{0}$
@@
We next adapt example 1 to illustrate our two leading cases, in the context of engaging test-motivated students. Suppose the true state of the world $T$ consists of both a test-relevant and test-irrelevant dimension, $T=T_{1}\times T_{2}=\{L,R\}\times\{0,1\}$. The student and teacher know that the student will be asked a single question, $Q=\{Q_{0}\}$, whose responses are $A=\{l,r\}$, with $Q_{0}(L0\}=Q_{0}(L1)=l$ and $Q_{0}(R0)=Q_{0}(R1)=r$. That is, the $\{L,R\}$ component of the true state is relevant, and the $\{0,1\}$ component is irrelevant. The student (agent) again faces a binary choice $A=\{l,r\}.$ The student’s utility is one if she answers the question correctly and negative one otherwise, and therefore is identical (ignoring the test-irrelevant dimension) to that in example 1. The prior belief is uniform. We continue to assume the agent’s information processing capacity is defined by the negative of Shannon’s entropy, and continue to use the parameters $\kappa=2,\chi=1$.
@@
We first study the leading case of $H=G$, and illustrate Proposition 4 with a specific example. Like Example 1, both the principal’s and the agent’s optimization problems are equivalent to static rational inattention problems. By re-parameterizing the problems using the condition probability of the chosen action, both problems can be rewritten as:
$\sup_{P(a|t_{1},t_{2})}\sum_{a,t_{1},t_{2}}\frac{1}{4}u(a,t_{1})P(a|t_{1},t_{2})-C\cdot\bigg($ $\sum_{a}\sum_{t_{1},t_{2}}\frac{1}{4}P(a|t_{1},t_{2})\log(P(a|t_{1},t_{2})$
$-\sum_{a}\bigg(\sum_{t_{1},t_{2}}\frac{1}{4}P(a|t_{1},t_{2})\bigg)\log\bigg(\sum_{t_{1},t_{2}}\frac{1}{4}P(a|t_{1},t_{2})\bigg)\bigg),$
for $C=2$ (the agent-optimal problem) and $C=2-\lambda$ (the principal-optimal problem). Note that replacing $P(a|t_{1},t_{2}=0)$ and $P(a|t_{1},t_{2}=1)$ with $\frac{P(a|t_{1},0)+P(a|t_{1},1)}{2}$ (denoted by $P(a|t_{1})$) does not change the positive term while strictly reduces the negative term if $P(a|t_{1},0)$ and $P(a|t_{2},1)$ are not identical. Evidently, w.l.o.g., the optimization problem reduces to
$\sup_{P(a|t_{1})}\sum_{a,t_{1}}\frac{1}{2}u(a,t_{1})P(a|t_{1})-C\cdot\bigg($ $\sum_{a,t_{1}}\frac{1}{2}P(a|t_{1})\log(P(a|t_{1}))$
$-\sum_{a,t_{1}}\frac{1}{2}P(a|t_{1})\log\bigg(\sum_{t_{1}}\frac{1}{2}P(a|t_{1})\bigg)\bigg),$
which is equivalent to a static rational inattention problem with $t_{1}$ being the only unknown state. [^20] This equivalence is special to mutual information and is the implication of a more general “compression invariance” property introduced by caplin2017shannon; bloedel2020cost. As shown in Proposition 4, when there is a test-irrelevant state that is costly to learn about, no information about the state will be acquired either in the student-optimal benchmark or in the principal-agent problem, even though the student learning about state $t_{2}$ enters the teacher’s payoff function.
@@
Now, we turn to the second leading case, in which the teacher only cares about how much the student knows about the state $t_{2}\in T_{2}$ (the test-irrelevant dimension). We assume that the teacher’s payoff is $2\mathbb{E^{\pi}}[|q_{t_{2}}-0.5|]$ if the measure of stopping belief is $\pi$, where $q_{t_{2}}$ denotes the probability of $t_{2}=1$ under the student’s stopping belief. Note that this payoff function is equivalent to the instrumental value of information in a hypothetical binary decision problem where the utility of matching (mismatching) the state $y$ is 1 (-$1$). Theorem 1 shows that the teacher’s optimal policy maximizes the following auxiliary problem (for some $\lambda>0$):
$\sup_{\pi}\mathbb{E^{\pi}}[2|q_{t_{2}}-0.5|+\lambda\widehat{u}(q)-2\lambda(H(q)-H(\bar{q}_{0}))],$
which is equivalent to a hypothetical static RI problem where the decision maker has four possible actions, whose utilities are
@@
Per matejka2015rational, the solution is logit, given by